At what distance from a source of light will the intensity of illumination be half as great as it is at a distance of 6 inches
Answer by Alan3354(69443)

At what distance from a source of light will the intensity of illumination be half as great as it is at a distance of 6 inches
------
I = k/d^2 --- k = a constant.
d^2 = k/I
-----------
d^2 = (k/(I/2)) = 2k/I
d = 6*sqrt(2) inches
----------------------------------------
That might be confusing.
---
Illumination intensity is a function of the inverse of the square of the distance.
d*2 = I/4
d*3 = I/9
d*x = I/x^2 --> d*sqrt(x) = I/x
-----------------
Inverse of the square of distance is common:
gravity, magnetism, E-field (of a point), sound intensity, and others.